### v1

# On Geometry-Enhanced Parameter-Efficient Fine-Tuning for 3D Scene Segmentation

Abstract
The emergence of large-scale pre-trained point cloud models has significantly advanced 3D scene understanding, but adapting these models to specific downstream tasks typically demands full fine-tuning, incurring high computational and storage costs. Parameter-efficient fine-tuning (PEFT) techniques, successful in natural language processing and 2D vision tasks, would underperform when naively applied to 3D point cloud models due to significant geometric and spatial distribution shifts. Existing PEFT methods commonly treat points as orderless tokens, neglecting important local spatial structures and global geometric contexts in 3D modeling. To bridge this gap, we introduce the Geometric Encoding Mixer (GEM), a novel geometry-aware PEFT module specifically designed for 3D point cloud transformers. GEM explicitly integrates fine-grained local positional encodings with a lightweight latent attention mechanism to capture comprehensive global context, thereby effectively addressing the spatial and geometric distribution mismatch. Extensive experiments demonstrate that GEM achieves performance comparable to or sometimes even exceeding full fine-tuning, while only updating 1.6% of the model’s parameters, fewer than other PEFT methods. With significantly reduced training time and memory requirements, our approach thus sets a new benchmark for efficient, scalable, and geometry-aware fine-tuning of large-scale 3D point cloud models. Code will be released.
Figure 1: (a) Existing PEFT methods, such as adapters, prompt tuning, and LoRA, focus on adaptations inside attention and feed-forward layers. (b) In contrast, Geometry Encoding Mixer (GEM) explicitly encodes the geometric cues and mixes them into the pre-trained model, by the Spatial Adapter refining the pre-trained positional encoding and the Context Adapter complementing the local attention. (c) By capturing 3D spatial details and scene-wide geometry context, GEM reaches full fine-tuning performance while tuning $<2\%$ parameters, outperforming existing PEFT methods.
1 Introduction
Point cloud semantic segmentation is a fundamental task for scene understanding that underpins many real-world applications, from autonomous driving and unmanned aerial vehicles to augmented reality [20, 81, 46]. Recent breakthroughs in large-scale models for language and 2D vision [16, 50, 58, 14, 57, 49] have spurred an analogous trend toward training powerful point cloud backbones capable of capturing rich semantics from 3D scenes [74, 75, 77].
Despite significant capabilities, adapting large pre-trained backbone models to specific downstream target domains is non-trivial. The typical strategy heavily relies on full fine-tuning, a process that is both computationally expensive and storage-intensive, since each downstream task demands an independent copy of all model parameters. Furthermore, fine-tuning large models typically requires extensive datasets to prevent overfitting and catastrophic forgetting [3, 51], while large datasets are usually inaccessible when adapted to smaller, specialized tasks. To address these issues, parameter-efficient fine-tuning (PEFT) methods offer a promising performance in reducing the cost of adaptation while reaching the performance of full fine-tuning with significantly fewer parameters.
While PEFT has been extensively validated in language processing and 2D vision tasks, its effectiveness on large-scale 3D point clouds remains largely unexplored and inadequately addressed. As demonstrated empirically in Fig. 1, existing PEFT methods only yield limited performance gains if directly applying them to a 3D point cloud backbone. We tend to attribute this limitation to a key challenge that is often overlooked by current PEFT methods when using for 3D point cloud: unlike structured data such as text or images, point clouds are inherently unordered sets of coordinates in $\mathbb{R}^{3}$. They exhibit strong irregularity, sparsity, and structural variability, shaped by different sensing protocols and scene geometries. These factors lead to significant geometric and spatial distribution shifts between large-scale pre-training datasets and downstream domains, which existing PEFT methods fail to account for. In particular, representative PEFT methods, such as LoRA [26], adapters [25] and prompt tuning [37, 36], either adapt models at an isolated, per-point level or insert fixed external tokens at the global level, thus failing to adequately capture important geometric and spatial contexts inherent in 3D scenes. Moreover, to avoid the prohibitive computational cost of global attention over a large number of points, current 3D transformers predominantly employ local attention mechanisms without explicitly modeling global contexts, which further restricts the potential of current PEFT approaches in performance.
To address the above challenges, we re-examine PEFT for 3D scene understanding and hypothesize that effective PEFT on 3D scenes could take advantage of explicitly modeling both fine-grained local spatial patterns and global geometric contexts. We propose that neglecting either aspect would degrade fine-tuning performance: local-only adaptations may be prone to noise due to the lack of broader context consideration, whereas global-only methods would miss local details and lose precision. This motivates a novel PEFT adaptation framework tailored specifically for 3D scenes to bridge these local and global scales.
We introduce Geometry Encoding Mixer (GEM), a geometry-aware module for parameter-efficient fine-tuning of point cloud transformer on 3D scenes. It comprises two complementary components: a spatial adapter at local neighborhood and a context adapter capturing the scene context. In the spatial adapter, we employ a lightweight 3D convolutional bottleneck that operates on points neighborhood, enriching the pre-trained positional encoding by learning fine-grained local spatial details at target domains. In the context adapter, we introduce a set of learned latent tokens to serve as global context vectors. These latent tokens interact with the full point cloud through efficient attention, forming a bottleneck at the token dimension and bypassing the constraints of local attention, and thus aggregating scene-specific context from across the entire point cloud. By fusing these two paths, the Geometry Encoding Mixer effectively bridges local and global representations, providing the fine-tuned model with a richer understanding of the 3D scene than either component alone.
Empirically, we validate our approach on large-scale 3D scene datasets, including both indoor [12, 60, 1, 82] and outdoor [4] scenes. Our results indicate that models equipped with GEM consistently achieve performance matching or sometimes surpassing full fine-tuning methods, updating merely ~1.6% of model parameters, while being fewer than existing PEFT alternatives. These findings underscore the importance of explicitly modeling geometric and spatial contexts for efficient and effective adaptation in 3D scene understanding. To the best of our knowledge, our work represents the first exploration and validation of PEFT approaches tailored explicitly for large point cloud transformer under large-scale 3D scenes, hoping to establish a foundation for future research and practical deployments.
2 Related Work
Point cloud segmentation.
Point clouds serve as an efficient representation for large-scale 3D scenes. Consequently, point cloud segmentation becomes a fundamental task that drives the design of 3D backbone architectures. Since the seminal work PointNet [53, 54], numerous 3D backbone architectures have been proposed for this task. These backbones either project the points onto a grid-like structure, e.g. 3D voxel grid, to exploit 3D convolutional networks [10, 69, 64, 21, 19, 13], or directly process the raw unordered point sets [71, 38, 73, 27, 65, 45]. Recently, inspired by the success of large transformer models in natural language and 2D vision [67, 16, 6, 9], researchers have started training increasingly large and powerful point cloud backbones [55, 88, 68, 56, 34, 76, 75]. Although these models achieve strong performance, they typically need to be trained from scratch for each new dataset and often require specialized training recipes to reach optimal results.
Following the paradigm of large-scale pre-training that has proven successful in vision and language domains [58, 33, 50, 14, 49, 23], researchers have begun exploring similar strategies for 3D point clouds. In particular, self-supervised learning (SSL) on large 3D scene datasets has demonstrated promising results [80, 44, 24, 78]. However, most such efforts focus on modest convolutional backbones like SparseUNet [10] rather than transformer-based architectures. Only recently has an SSL approach been applied to a transformer-based 3D backbone: Sonata [74] introduced SSL for a large point transformer encoder [75], but even this method still requires full fine-tuning on each downstream dataset to achieve optimal performance.
In this work, we explore effective PEFT methods for large pre-trained point cloud backbones, with the goal of reducing the computational and storage overhead when adapting them to downstream datasets, while matching the performance of full fine-tuning.
Parameter-efficient fine-tuning (PEFT).
The ever-growing size of transformer-based foundation models[16, 33, 66] makes full fine-tuning for downstream task prohibitively expensive in both memory and computation. PEFT methods address this challenge by adapting large models while updating only a small fraction of their parameters. Existing approaches fall into four broad categories.
Selective fine-tuning methods update a carefully chosen subset of the original weights. One simple variant is linear probing [50, 15] that trains only the classification head. Other strategies restrict training to specific parts of the network, e.g., tuning only bias terms [5] and selecting a subset of parameters based on gradient magnitude criteria [87].
Adapter-based tuning inserts lightweight bottleneck modules into an otherwise frozen backbone. First explored for CNNs [59] and later extended to transformers, adapters may be placed sequentially [25, 61] or in parallel [22, 90, 17, 8] to the original modules. Beyond single task, adapters can be composed or fused, promoting knowledge sharing without catastrophic forgetting [70, 51].
Prompt-based tuning techniques, including prompt tuning [36] and prefix tuning [37], extend to the prompting paradigm in large language models [41]. Instead of hand-crafted prompts, these methods learn task-specific vectors that are prepended to the input sequence [2, 36] or injected as extra tokens in transformer layers [37, 43, 31], thereby steering the model without modifying its original weights [36, 41].
Low-rank adaptation (LoRA) constrains the weight updates to a low-dimensional subspace, by learning a pair of low-rank matrices that are added to each pre-trained weight, typically in the attention layers [26]. This design approximates the effect of full fine-tuning while introducing only a small number of additional parameters. Successors further enhance the approximation ability, robustness, or rank allocation [32, 42, 84, 79, 40].
Recent works also propose hybrid schemes that highlight common design principles [28], for instance, integrating their respective strengths under a unified adapter framework [22].
Although PEFT has proved effective in language and 2D vision, a direct transfer to 3D scene understanding is often sub-optimal, largely due to the unordered nature of point clouds and the prominence of geometric cues. Our work thus investigates PEFT strategies tailored to large-scale 3D scenes.
PEFT for 3D point cloud.
In the context of 3D point clouds, PEFT remains relatively underexplored. Existing methods thus far have focused mostly on object-level inputs with limited spatial scale. Some approaches introduce prompt tokens that adapt to 3D data, where the prefix tokens are dynamically generated from intermediate features [83, 89] or from the spatial centers of local patches [85]. Several works also introduce auxiliary graph-based side networks, operating either in the spectral domain [39] or the spatial domain [18], to provide global geometric context. Other methods construct banks of prompt vectors using domain-specific dataset to inject 3D prior knowledge [63], or combine language-side prompt tuning with point-cloud adapters to handle open-vocabulary recognition [62].
Compared to object-level datasets [7, 48], 3D scene understanding involves much larger inputs containing on the order of millions of points [12, 1]. This scale amplifies computational challenges and underscores the need for PEFT techniques expressly designed for large 3D scene models, which is however a research direction that remains largely unexplored.
3 Methodology
We propose the Geometry Encoding Mixer (GEM) as a lightweight parameter-efficient module for fine-tuning point-cloud transformer. Fig. 2 summarises the overall architecture.
In particular, GEM consists of two complementary adaptations: a Spatial Adapter that refine the positional encoding of each point, injecting local geometry information overlooked by generic adapters; and a Context Adapter that distills a compact set of latent tokens that broadcast global scene-level cues. Both components follow the residual, bottleneck design of classic adapters, yet emphasize and operate on spatial rather than channel space.
3.1 Preliminaries
We first consider the direct application of existing PEFT methods on current point-cloud transformers.
Chanlleges under 3D.
Following transformer [67, 16], point-cloud transformers also build upon self-attention layer (Attn) followed by feed-forward network (FFN). However, due to the large scale input points at million scale, global self-attention over all points would demand prohibitively large GPU memory and compute resource, if ever possible, due to the quadratic complexity of the attention operations: $\mathcal{O}(\text{Attn})=n^{2}$, where $n$ denotes the number of input points. In this regards, state-of-the-art point-cloud transformers propose various design of local attentions [88, 76, 75, 34, 68] to cap the complexity, where each point can only attend to points within the same patch. With a patch size of $p$, the complexity of local attention is then $\mathcal{O}(\text{Attn}_{\text{loc}})=np$. For example, it is set to $p=16$ in PT [88] and $p=1024$ in the larger PTv3 [75].
The irregular and sparse nature of 3D point cloud, together with these architecture design, greatly imped the application of standard PEFT methods such as adapter, LoRA and prompt tuning. As shown in Fig. 1, directly applying existing methods yields only marginal gains. We attribute this to their lack of spatial structure awareness of 3D geometry and the inability to capture scene-wide context under the local attention regimes of modern point cloud backbones.
Adapters.
The adapter-based methods [25, 61] insert small modules after the attention or FFN layers:
$f_{\text{adapter}}(\bm{x})=\bm{x}+\sigma(\bm{x}\bm{W}_{\text{down}})\bm{W}_{\text{up}},$ (1)
where $\bm{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{r\times d}$ are down and up projections that forms a bottleneck structure with $r\ll d$ and activation $\sigma$ in the mid, and are also surrounded by residual connection. As generic MLP bottlenecks, adapters adapt the model on a per-point basis, overlooking the important spatial cues in 3D point cloud.
LoRA.
LoRA methods [26] typically adapt the attention projection layers and approximate the full fine-tuning with a bottleneck in weights space, transforming the attention into:
$\text{Attn}_{\text{LoRA}}=\text{softmax}\frac{(\bm{Q}+\Delta\bm{Q})(\bm{K}+\Delta\bm{K})^{T}}{\sqrt{d_{k}}}\bm{V},$ (2)
where $\Delta\bm{Q}=\bm{x}\bm{W}_{\text{down}}\bm{W}_{\text{up}}$, with $\bm{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{r\times d}$, and similarly for $\Delta\bm{K}$ with a different pair of weights. While LoRA can leverage the strong ability of the attention layers, we note that all attentions in point cloud transformer are constrained within local patches. LoRA methods are thus inherently limited by the prevalent design of local attention. With only limited rank $r<<d$, it could fail to provide global context of the target dataset.
Pompt tuning.
Prompt tuning methods [37, 36] explicitly introduce global tokens into the attention:
$\text{Attn}_{\text{prompt}}=\text{softmax}\frac{\bm{Q}[\bm{P}_{K};\bm{K}]^{T}}{\sqrt{d_{k}}}[\bm{P}_{V};\bm{V}],$ (3)
where $\bm{P}_{K},\bm{P}_{V}\in\mathbb{R}^{m\times d}$ are two set of tokens prepending to the original key-value pair. In spite of the constraint of local patches, prompt tokens can be duplicated into each patch. However, with $m<<n$, it can be hard for a few static external tokens to capture scene-specific context that can vary across point clouds within the same dataset. In addition, it also overlooks the spatial pattern of the point cloud, lacking the adaptation on local point features.
Others.
There are more methods that explore other aspects. Selective tunings propose to tune a small subset of the forzen model parameters, such as BitFit [5] that tunes only bias terms. Hybrid methods are also explored, such as MAM [22] that couples FFN adapter with prompt tuning for attention layer. These methods yet still inherit the constraints from those representative PEFT methods, which are hindered by the local attention and fails to recognize the spatial structure.
In summary, existing methods overlook 3D geometry and remain confined to local patches, motivating a PEFT design that incorporates spatial awareness with scene-wide context, introduced next.
Figure 2: Geometry Encoding Mixed. We propose the spatial adapter to enhance the pre-trained positional encoding, and the context adapter to overcome the local attention mechanism, thus enhancing the efficient adaptation on large-scale 3D scenes with explicit geometry encoding.
3.2 Our Methodology: Geometry Encoding Mixer (GEM)
To address these limitations, we present the Geometry Encoding Mixer (GEM). It refines positional encodings and injects scene context for efficient adaptation on large-scale 3D scenes:
$\bm{x}$ $\leftarrow\bm{x}+\text{pos}(\bm{x})+f_{\text{spatial}}(\bm{x}),$ (4)
$\bm{x}$ $\leftarrow\bm{x}+\text{Attn}_{\text{loc}}(\bm{x})+f_{\text{context}}(\bm{x}),$ (5)
where $\bm{x}$ denotes input points and $\text{pos}(\cdot)$ is the pre-trained positional embedding. Fig. 2 illustrats the overall structure.
Spatial Adapter (SA).
Point clouds are sparse, irregular samples in 3D. Adapting to their fine-grained geometry requires explicitly modeling local neighborhoods, which prior PEFT methods fail to address.
To this end, the spatial adapter refines per-point positional encoding through a lightweight 3D convolutional bottleneck. Concretely, for each point $\bm{x}$, we consider a 3D grid and gather points in the vicinity voxels as neighbors, denoted by $\mathcal{N}$:
$f_{\text{spatial}}(\bm{x})=\bm{x}+\sigma(\sum_{i\in\mathcal{N}}\bm{W}^{i}_{\text{loc}}(\bm{x}\bm{W}_{\text{down}}))\bm{W}_{\text{up}},$ (6)
where $\bm{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{r\times d}$ are down and up projectsion with $r\ll d$. $\bm{W}^{i}_{\text{loc}}\in\mathbb{R}^{r\times r}$ composes the spatial kernel weights for $i$-th neighboring voxel and $\sigma(\cdot)$ is the nonlinearity, such as ReLU.
With a common kernel dimension $k=3$, the spatial adapter touches at most $k^{3}$ neighbours per point and adds $2rd+k^{3}r^{2}$ parameters, rendering a complexity of $\mathcal{O}(2ndr+nk^{3}r^{2})=\mathcal{O}(nd)$ given $k,r\ll n,d$. It thus functions as an efficient convolution-based positional encoding [11, 29] in parallel to the pre-trained positional encoding, thereby capturing fine spatial layout to match the target distribution.
Context Adapter (CA).
Due to the local attention, any local and channel-wise adaptation would be constrained from acquiring scene-wide context, hindering the adaptation performance on downstream tasks like semantic segmentation.
To inject global context without breaking the $\mathcal{O}(n^{2})$ barrier, we introduce $m$ latent tokens $\bm{L}\in\mathbb{R}^{m\times r}$, where $m\ll n$, that attend to all points once:
$\bm{L}_{c}$ $=\text{Attn}\bigl{(}\bm{L}\bm{W}^{Q},\,\bm{K},\,\bm{V}\bigr{)},$ (7)
$f_{\text{context}}(\bm{x})$ $=\bm{x}\;+\;\text{Attn}\bigl{(}\bm{Q},\,\bm{L}_{c}\bm{W}^{K},\,\bm{L}_{c}\bm{W}^{V}\bigr{)}\bm{W}_{\text{up}},$ (8)
where $\bm{L}\in\mathbb{R}^{m\times r}$ and $\bm{Q},\bm{K},\bm{V}=\bm{x}\bm{W}_{\text{down}}^{Q},\bm{x}\bm{W}_{\text{down}}^{K},\bm{x}\bm{W}_{\text{down}}^{V}$ are down-projected point features.
As both attention costs $O(nm)$ with $m\ll n$, global aggregation is affordable and at the same level of complexity as the conventional prompt tuning [37, 36]. Furthermore, we update $\bm{L}\leftarrow\bm{L}+\bm{L}_{c}$ at each adapter insertion to share across layers, yielding dynamic prompts that capture the context of current scene, unlike static prefixes in prompt tuning.
Discussion.
In comparison to existing PEFT methods, GEM explicitly biases the adaptation to 3D geometry and context. The spatial adapter captures localized variations that generic adapters, such as LoRA and basic adapters, would miss, since those ignore spatial structure and use the positional encodings from pre-training. The context adapter circumvent the local attention design to provide scene-wide context for efficient adaptation with few parameters. By combining these two components, our approach enables scene-aware adaptation of 3D transformers, where each point is tuned with respect to both its neighbors and the entire point cloud. It thus results in significantly improved performance on 3D scene datasets with only a lightweight adaptation overhead.
4 Experiments
We present the results of our proposed GEM on semantic segmentation of large-scale 3D scene datasets and experiment with both self-supervised pre-trained backbone, Sonata [74], and another one, PTv3-PPT [77], for investigation. We also provide ablation studies to better reveal the detailed effects of different components.
4.1 Experimental Setup
We primarily follow the comprehensive protocols proposed in Sonata [74], covering ScanNet [12], ScanNet200 [60], ScanNet++ [82] and S3DIS [1].
We adopt two representative pre-trained backbones, the Sonata model [74] with large-scale self-supervised training and the PTv3-PPT [77] with supervised pre-training on multiple large-scale curated datasets. We compare GEM with the existing popular PEFT methods, including bias-tuning from BitFit [5], Adapter [25], LoRA [26], and Prompt Tuning [37]. We consider linear probing as a baseline and the full fine-tuned model as our target strong reference. In addition, we note that the Sonata (full.) specifically introduce a task-specific decoder for semantic segmentation during the fine-tuning, which may not reflect the fair comparison. Therefore, we revive to the standard full fine-tuning without additional decoder network, which is denoted by Sonata (ft.).
For training, we follow the widely accepted fine-tuning setups to update only the inserted or selected weights with the pre-trained backbone weights remaining frozen. All PEFT baselines follow the implementations from released code, adopt the suggested common practice [47, 52], and are tuned to their best validation setting in Fig. 1(c). Specifically, We set the default rank to be $r=32$ and the number of learnable tokens to be $m=4$. More details are given in the supplementary.
4.2 Performance Comparison
Semantic Seg. Params ScanNet Val [12] ScanNet200 Val [60] ScanNet++ Val [82] S3DIS Area 5 [1] S3DIS 6-fold [1]
Methods Learn. Pct. mIoU mAcc allAcc mIoU mAcc allAcc mIoU mAcc allAcc mIoU mAcc allAcc mIoU mAcc allAcc
Training from scratch
SparseUNet [10] 39.2M 100% 72.3 80.2 90.0 25.0 32.9 80.4 28.8 38.4 80.1 66.3 72.5 89.8 72.4 80.9 89.9
PTv3 [75] 124.8M 100% 77.6 85.0 92.0 35.3 46.0 83.4 42.1 53.4 85.6 73.4 78.9 91.7 77.7 85.3 91.5
Full fine-tuning
Sonata (full.) [74] 124.8M 100% 79.4 86.1 92.5 36.8 46.5 84.4 43.7 55.8 86.6 76.0 81.6 93.0 82.3 89.9 93.3
Sonata (ft.) 108.5M 100% 78.3 85.9 92.3 37.3 47.8 83.7 49.8 61.2 87.6 72.4 79.0 92.2 79.5 87.3 92.3
PEFT methods
Sonata (lin.) 0.02M 0.02% 72.5 83.1 89.7 29.3 41.6 81.2 37.3 50.9 84.3 72.3 81.2 90.9 76.5 87.4 90.8
+ BitFit [5] 0.2M 0.2% 74.7 84.7 90.8 32.5 45.5 82.0 42.4 56.5 85.7 73.9 82.0 91.5 76.1 87.1 91.0
+ Adapter [25] 2.8M 2.5% 77.0 85.4 91.8 33.6 45.7 82.5 42.6 57.5 85.6 73.8 82.9 91.5 76.4 87.5 91.4
+ LoRA [26] 1.9M 1.7% 76.7 85.7 91.7 33.6 45.5 82.7 44.2 58.0 86.5 74.5 83.2 91.5 77.4 87.8 91.2
+ Prompt Tunning [37] 5.5M 4.8% 74.3 84.1 90.5 31.4 44.4 81.6 41.2 56.2 84.8 73.4 82.5 91.0 73.7 86.5 90.5
+ GEM (ours) 1.8M 1.6% 78.3 86.6 92.3 35.6 46.9 83.3 46.6 60.3 86.3 75.1 83.0 92.1 77.9 88.2 92.1
Table 1: Semantic segmentation.
Main results.
Tab. 1 shows that GEM consistently surpasses all representative PEFT methods across datasets. It matches the performance of full fine-tuning on most datasets and even exceeds it on ScanNet++[82]. ScanNet++ comprises large and diverse scenes captured at sub-millimeter resolution, diverging greatly from the spatial distribution of common pre-training datasets such as ScanNet [12]. Under this domain gap, the joint modeling of local spatial cues and global scene context introduced by GEM proves especially beneficial.
Interestingly, LoRA [26] and adapter [25] deliver similar scores, despite acting on different transformer components. This observation implies that, when updates remains local and geometric is not modeled explicitly, the precise choice of adaptation target can be secondary. In addition, prompt tuning [37] underperforms even the linear probing baseline in the S3DIS 6-fold evaluation, revealing the penalty of ignoring spatial structure during fine-tuning.
Data efficiency.
We assess PEFT methods under the scenarios of limited data and annotations in Tab. 3. The results demonstrate the exceptional data efficiency of GEM, which outperforms both other PEFT methods and the full fine-tuning counterparts. Notably, under extreme data scarcity, such as 1% of the labeled scenes and limited annotations (20 points per scene), GEM surpasses both Sonata (full.) and Sonata (ft.), highlighting its superiority in low-data regimes.
Supervised pre-training.
While self-supervised methods dominate in the language domain, supervised pre-training remains competitive for vision [33]. To probe the limits of PEFT under large-scale 3D supervision, we adopt recent advances in 3D supervised pre-training [77], which achieve leading performance by training on a large collection of curated, labeled scene datasets. As shown in Tab. 3, existing PEFT methods can even degrade performance, likely suffering from negative transfer [86, 72]. In contrast, GEM improves upon supervised backbone, further outperforming the larger PTv3 [75] with only 1.6% additional parameters.
PEFT with decoder.
A dedicated segmentation decoder ( 13% of the total parameters) is known to lift fully fine-tuned baselines. We therefore ask whether PEFT can still add value in the presence of such a task-specific head.
Table 5 answers in the affirmative: GEM outperforms all competing PEFT variants and even surpasses its own fully fine-tuned counterpart, setting a new state-of-the-art on ScanNet. The result underscores the merit of jointly modeling local geometry and scene-level context, even when the decoder already provides task-specific capacity.
Outdoor segmentations.
During the original evaluation of Sonata [74], it employs separate pre-trained models for indoor and outdoor scenarios, due to the significant domain gaps between indoor and outdoor scenes. Here we test a stronger setting: transferring an indoor pre-trained backbone to an outdoor autonomous-driving benchmark. We replace the input layer with a randomly initialized counterpart to match dimensionality and freeze all remaining weights.
Table 5 shows that GEM narrows the gap to a model trained from scratch on outdoor data, yet a noticeable performance gap persists. Closing this gap remains an interesting direction for future work.
[…]
Data Efficiency Limited Scenes (Pct.) Limited Annotation (Pts.)
Methods 1% 5% 10% 20% Full  20 50 100 200 Full
Training from scratch
SparseUNet [10] 26.0 47.8 56.7 62.9 72.2 41.9 53.9 62.2 65.5 72.2
PTv3 [75] 25.8 48.9 61.0 67.0 77.2 60.1 67.9 71.4 72.7 77.2
Full fine-tuning
Sonata (full) [74] 45.3 65.7 72.4 72.8 79.4  70.5 73.6 76.0 77.0 79.4
Sonata (ft.) 44.4 63.2 71.3 72.3 78.3 69.6 72.6 75.3 76.2 78.3
PEFT methods
Sonata (lin.) 43.6 62.5 68.6 69.8 72.5 69.0 70.5 71.1 71.5 72.5
+ BitFit [5] 46.5 64.9 69.9 71.7 74.7 71.0 72.3 72.9 73.6 74.7
+ Adapter [25] 46.4 64.2 70.1 72.2 77.0 71.8 73.4 74.7 75.0 77.0
+ LoRA [26] 46.6 63.0 70.1 72.6 76.7 72.1 73.6 75.2 75.5 76.7
+ Prompt Tunning [37] 45.5 62.6 68.9 71.1 74.3 70.2 71.5 72.5 72.8 74.3
+ GEM (ours) 47.5 65.6 71.0 73.3 78.3 72.3 74.7 76.2 76.6 78.3
Table 2: Data efficiency.
Supervised Pre-train. Params ScanNet Val [12]
Methods Learn. Pct. mIoU mAcc allAcc
Training from scratch
SparseUNet [10] 39.2M 100% 72.3 80.2 90.0
PTv3 [75] 124.8M 100% 77.6 85.0 92.0
Full fine-tuning
PTv3-PPT (ft.) [77] 97.4M 100% 78.6 86.0 92.5
PEFT methods
PTv3-PPT (lin.) 0.1M 0.1% 78.6 85.9 92.5
+ BitFit [5] 0.2M 0.2% 78.2 85.6 92.3
+ Adapter [25] 1.8M 1.8% 78.5 85.9 92.4
+ LoRA [26] 1.4M 1.7% 78.4 86.0 92.4
+ Prompt Tunning [37] 4.8M 4.8% 78.3 85.9 92.4
+ GEM (ours) 1.8M 1.6% 79.1 86.6 92.6
Table 3: Supervised pre-training.
[…]
with Decoder Params ScanNet Val [12]
Methods Learn. Pct. mIoU mAcc allAcc
Training from scratch
PTv3 [75] 124.8M 100% 77.6 85.0 92.0
Full fine-tuning
Sonata (full.) [74] 124.8M 100% 79.4 86.1 92.5
Sonata (full.) 124.8M 100% 78.5 86.3 92.4
PEFT methods with decoder
Sonata (dec.)  16.3M 13.1% 79.1 86.6 92.7
Sonata (dec.) 16.3M 13.1% 77.2 85.9 91.9
+ BitFit [5] 16.4M 13.2% 78.1 86.1 92.3
+ Adapter [25] 19.1M 15.0% 78.2 86.5 92.3
+ LoRA [26] 18.2M 14.3% 78.9 87.0 92.5
+ Prompt Tunning [37] 22.6M 17.2% 77.4 86.1 92.1
+ GEM (ours) 18.1M 14.3% 79.5 87.3 92.7
Table 4: PEFT with segmentation decoder. Methods with reports our re-produced results.
Outdoor Seg. Params Sem.KITTI Val [4]
Methods Learn. Pct. mIoU mAcc allAcc
Training from scratch
PTv3 [75] 124.8M 100% 69.1 76.1 92.6
Full fine-tuning
Sonata (full.) [74] 124.8M 100% 72.6 77.9 93.4
Sonata (ft.) 108.5M 100% 68.8 75.2 92.6
PEFT methods
Sonata (lin.) 0.02M 0.02% 62.0 72.5 91.0
Sonata (lin.) 0.02M 0.02% 52.0 63.3 87.1
+ BitFit [5] 0.2M 0.2% 59.9 70.6 91.0
+ Adapter [25] 2.8M 2.5% 62.5 72.2 92.0
+ LoRA [26] 1.9M 1.7% 63.8 74.0 92.0
+ Prompt Tunning [37] 5.5M 4.8% 59.1 71.0 90.8
+ GEM (ours) 1.8M 1.6% 67.7 75.5 93.1
Table 5: Outdoor semantic segmentation. Methods with use outdoor datasets for pre-training.
4.3 Ablations and Analsys
We mainly consider the linear probing, Sonata (lin.) as baseline and ablate on ScanNet [12] to better investigate the key factprs for effective PEFT in 3D scenes, shown in Tab. 6. For more studies on the effect of available parameters, please refer to the supplement.
Individual effectiveness of SA and CA.
To validate our initial hypothesis, we study the individual effectiveness of the SA and CA in LABEL:tbl:abl_comp. We find that both adapters can can already be superior to the baseline and competitive to the existing PEFT methods. We notice that the SA takes the majority part of the introduced paramteres, which is partially due to the inherently expensive 3D convolutoin with $k^{3}$ kenrels. Such challenge is also 3D specific and also urges us to develop the CA, rather than soly relying on the positional encoding to capture 3D geometry. By combining the two form of spatial adapters, we obtain GEM that reaches the best performance by comprehensively exploring the 3D scene while remaining parameter-efficient.
Understanding the latent tokens.
In LABEL:tbl:abl_tokens, we are motivated to further investigate how efficient and effective of the proposed CA could be. We thus study how few tokens it could afford before failing in capturing the global context and obtaining a similar performance to SA-only. Surprisingly, we find that CA can use as few as one single token to effectively express the global scene context, and obtaining clear improvement over SA-only approach. We consider this to be the effectiveness of weight decomposition, where we are akin to approximating the $n^{2}$ global self-attention weights with two $n\times 1$ matrix. In comparison to LoRA that decompose the in weight space, we decompose the spatial attention weight matrices and form a spatial bottleneck at token dimension. In addition, while we are aware of such similar topic has been similarly explored for efficient attentions [30, 35], we are, to the best of our knowledge, the first to motivate their effective application as PEFT methods for 3D scenes understanding.
Sharing the latent tokens.
The last experiment, LABEL:tbl:abl_dist probes whether the latent tokens learned by CA should remain layer-local or be shared across the hierarchy, assessing its ability of capturing dynamic and scene-specific context. When each layer keeps its own bank of tokens (N/A), GEM already surpasses SA-only, confirming that CA supplies complementary global cues. Re-using the same tokens within each encoder stage (per-stage) yields a further yet modest gain, suggesting that a coherent context inside each resolution level helps the network reconcile local geometry with mid-range structure.
The strongest result emerges when a single set of latent tokens are shared by all transformer blocks. We attribute this improvement to its coincidence with the fact that 3D point cloud represent the same underlying 3D geometry in sipte of its irregularity and sparsity. From this perspective, such globally shared latent tokens confines the model to adapt to the true underlying scene, rather than overfitting on irrelavent pattern due to noisy sampled point cloud.
Revealing the geometry cues in latent tokens.
With the above assumption on latent tokens regularize the model to capture the underlying 3D scene, we visualize its attentional weights and compare to other similar PEFT method that can also inject global context, such as the prompt tuning.
As shown in Fig. 3, starting from the first stage, our latent tokens learn to produce much crisper attention weights following the geometric cues, such as the simple geometric primitives of floor surface. In deeper stages, latent tokens start to pick up more salient geometric objects, such as clutter on the table and sofa. In contrast, the static tokens introduced by prompt tuning produce attention blurry weights that span accross the boundaries of objects and scene areas. As a result, it is suggested that the proposed latent tokens can indeed deliver scene-wide context that broadcast the geometry cues. Together with the local spatial structure extracted by 3D convolutions, it thus results in its superior performance as an effective PEFT methods for 3D scenes.
SA  CA Params  mIoU
Sonata (lin.) 0.02% 72.5
+ GEM 1.0% 77.2
0.6% 77.3
1.6% 78.3
$m$ mIoU mAcc allAcc
1 78.1 86.4 92.1
4  78.3  86.6  92.3
8 78.2 86.8 92.3
  strategy mIoU mAcc allAcc
N/A. 77.8 86.1 92.1
per-stage. 78.0 86.1 92.0
global 78.3 86.6 92.3
Table 6: Ablations on GEM. If not specified, the backbone is the self-supervised pre-trained Sonata [74] with no additional task-specific decoder, evaluated on ScanNet [12]. Default settings are marked in gray.
Figure 3: We visualize the attention weights of our latent tokens, comparing to the attentional weights with the prompt tuning, showing the enhanced geometry cues produced by GEM.
5 Conclusion
This study presents GEM, a geometry-aware fine-tuning module tailored for large-scale 3D scene segmentation. By combining local spatial refinement and global context modeling, GEM addresses key challenges inherent in adapting pre-trained 3D models to new domains. Experimental results across multiple benchmarks confirm its ability to achieve competitive performance with minimal parameter updates. These findings highlight the value of incorporating geometric priors into efficient model adaptation and offer a promising direction for scalable deployment in 3D scene understanding tasks.
Limitation and future work.
While GEM achieve promising performance with little overhead in complexity, we notice it relies on the assumptaion that the model can be fully parallelized. For example, it would incur noticeable overhead if using large batch sizes, as the device is forced to process the input sequentially, in sipte of the parallel design in GEM. In addition, we are able to successfully train our PEFT methods with a fractional of epoch than the full fine-tining, e.g. 100 ranther than 800 epochs, we realize that the gradients need to be back-propagated into early layer and could hinder the training efficiency, especially with dense point cloud like S3DIS [1]. We thus suggest that it is required a systematic examination on the additional overhead that PEFT methods would experience in 3D scene understanding.
Supplementary
In this supplementary, we provide more materials as follow,
Appendix A  details the implementation, training, and inference;
Appendix B  presents additional experiments and analyses with tight parameter budgets
Appendix C  offers more qualitative visualizations.
Appendix A Details of Implementation, Training and Inference
Our implementation is based on the open-source codebase Pointcept (here) and follows the official implementations for Sonata [74], PPT [77] as well as PTv3 [75].
Leveraging the parameter efficiency of our method, we train on a single 4090 GPU for much fewer epochs to obtain the reported performance. For example, we train on ScanNet for 100 epochs, in contrast to the 800 epochs in the released Sonata configuration. Our code will also be released.
Table LABEL:tbl:profiling summarizes empirical latency and memory usage. We intentionally refrain from deployment-specific optimizations such as LoRA weight merging. Although the global attention and local convolution modules introduce extra cost, the runtime and memory footprint stay within the same order of magnitude as other PEFT baselines.
Appendix B More Experiments and Analysis
In spite of the promising results shown in Sec. 3, we suggest that it can better demonstrates the full potentials of PEFT methods under more challenging settings.
To stress parameter efficiency, we constrain all methods to the same limited learnable-parameter budgets, including enforcing rank $r=1$, allowing 0.1% parameters, and 1% parameters. As reported in LABEL:tbl:semseg_tight, GEM delivers consistently higher accuracy, whereas several baselines fail when the budget becomes extremely stringent.
Appendix C More Visualizations
We provide more qualitative results regarding the attention map generated by our latent tokens in LABEL:fig:attn-more. Despite shared across the stages, the latent tokens appear to focus on differen parts of the scenes, providing different scene context to the backbone models at different stages.